OpenStudy (anonymous):

Find the limit: Lim (x^3)-9x/x-3 x->3

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3}x ^{3}-9x \div x-3\]

OpenStudy (anonymous):

\[\frac{x^3-9x}{x-3}=\frac{x(x^2-9)}{x-3}=\frac{x(x+3)(x-3)}{x-3}=x(x+3)\]

OpenStudy (anonymous):

replace \(x\) by \(3\) for your answer

OpenStudy (anonymous):

Thank you man.